Attiya10:

Show that Z5×Z3 is cyclic group

Narad:

\[ℤ_{m} * ℤ _{n}\] is a cyclic group iff GCD(m,n)=1 that is m and n are relatively prime.

jhonyy9:

so like a first step hope you know what mean ,,cyclic group" and what are (mean) Z5 and Z3 ?

OronSH:

\(\mathbb Z_3\times\mathbb Z_5\) is isomorphic to \(C_{15}\)